OpenStudy (lgbasallote):

is it possible to use Laplace transformations on first order differential equations?

OpenStudy (anonymous):

it is

OpenStudy (lgbasallote):

could you show me a demonstration? for example... (x^2-y)y' + 2x^3 +2xy = 0

OpenStudy (anonymous):

it is not

OpenStudy (lgbasallote):

uhh so which is it?

OpenStudy (anonymous):

emm...can we evaluate L{yy'} ?

OpenStudy (lgbasallote):

i have no idea...can we?

OpenStudy (anonymous):

we cant

OpenStudy (lgbasallote):

aww so laplace transform is only possible in higher order diff eqs?

OpenStudy (anonymous):

not necessarily

OpenStudy (zzr0ck3r):

well its not the order here, is it? we could still use LT on y` = cos(x). this is 1st order

OpenStudy (zzr0ck3r):

its not the order that is the problem*

OpenStudy (lgbasallote):

you can do laplace on things with x? i thought you can only use it on t

OpenStudy (zzr0ck3r):

where x = t

OpenStudy (lgbasallote):

so laplace of that is sy(s) - y(0) = s/(s^2 + k^2) ??

OpenStudy (zzr0ck3r):

looks right to me

OpenStudy (zzr0ck3r):

yes had to check the old LT notes.

OpenStudy (lgbasallote):

so what happens after that?

OpenStudy (zzr0ck3r):

wait it should be y(t)

OpenStudy (zzr0ck3r):

then you would solve for y(t) and then try and find some inverse from a table of inverses that would help you solve

OpenStudy (zzr0ck3r):

I think.. its been a bit. one sec

OpenStudy (zzr0ck3r):

s*L(y(t)) - y(0) = s/(s^2 + 1^2)

OpenStudy (zzr0ck3r):

solve for L(y(t)) then take the inverse of both sides. no we will have some y(t) = L^(-1)(...) then you use a table of iverses. Most of the time you need ot do some algebra magic to get it to work out.

OpenStudy (lgbasallote):

lol. i guess sticking to the traditional way is better?

OpenStudy (zzr0ck3r):

well, sometimes not. most of the time yes.

OpenStudy (zzr0ck3r):

thought I should do it since I could not remember how

OpenStudy (zzr0ck3r):

its good for dirac delta function, stuff like that